# Physically Guided Deep Unsupervised Inversion for 1D Magnetotelluric Models

### v2

@@ Abstract
The global demand for unconventional energy sources such as geothermal energy and white hydrogen requires new exploration techniques for precise subsurface structure characterization and potential reservoir identification. Magnetotelluric (MT) method is crucial for these tasks, providing critical information on the distribution of subsurface electrical resistivity at depths ranging from hundreds to thousands of meters. However, traditional iterative algorithm-based inversion methods require the adjustment of multiple parameters, demanding time-consuming and exhaustive tuning processes to achieve proper cost function minimization. Although recent advances have incorporated deep learning algorithms for MT inversion, which have been primarily based on supervised learning, and needs large labeled datasets for training. This work utilizes TensorFlow operations to create a differentiable forward MT operator, leveraging its automatic differentiation capability. Moreover, instead of solving for the subsurface model directly, as classical algorithms perform, this paper presents a new deep unsupervised inversion algorithm guided by physics to estimate 1D MT models. Instead of using datasets with the observed data and their respective model as labels during training, our method employs a differentiable modeling operator that physically guides the cost function minimization, making the proposed method solely dependent on observed data. Therefore, the optimization algorithm updates the network weights to minimize the data misfit. We test the proposed method with field and synthetic data at different acquisition frequencies, demonstrating that the resistivity models obtained are more accurate than those calculated using other techniques. Our implementation is available at https://github.com/PAULGOYES/MT_guided1DInversion.git.
I Introduction
@@ III-A Experiment I:
We evaluated the proposed method on a simple three-layer model with theoretical thicknesses of 2500 m for each layer and the depth to the top of the half-space layer at 5000 m. The MT acquisition was simulated with a frequency range from 10-3 Hz to 102 Hz. For inversion, we used a ten-layer model with a thickness of 555 m for each layer. In this experiment, we compared the proposed method with the solution of (6) i.e., –norm, which was initialized with a 100 m constant model. Note that for both cases, we used the same forward MT operator. In this experiment we used a standard deviation
Fig. 2 shows the convergence curve for the cost function minimization. The –norm exhibits a smooth curve stabilizing after iteration 200 and reaching a minimum error of approximately 0.35. On the other hand, the proposed method, which includes the inverse neural operator, displays a steep curve in the first five iterations and stabilizes after iteration 125 with small fluctuations around approximately 0.15. This indicates that the proposed method achieves faster convergence of the data misfit for the inverse problem. Although both curves converged, Fig. 3a shows that the –norm formulation exhibits a mismatch, particularly at the lowest frequencies, whereas the proposed method fits all observations across all frequencies. This translates to differences in the inverted models in Fig. 3b, mainly in the last layer (i.e., at a depth of 5 km), where the proposed method shows a good fit. The computation times are 85.543 s and 81.346 s for the proposed and the –norm formulation, respectively.
@@ III-B Experiment II:
We tested the inversion algorithm’s capability on data where subsurface layers have variable thicknesses to simulate a realistic subsurface scenario. Our results were compared with the SimPEG inversion framework [9, 21], which incorporates modeling operators based on Maxwell’s solutions to simulate electromagnetic fields with plane waves and a Gauss-Newton approach that employs the conjugate gradient solver [22]. SimPEG was initialized with a homogeneous resistivity model of 100 m. For both inversion methods, we used 20 layers with the maximum depth set to 15,473 meters. The proposed method used epochs. In this experiment, we found the best parameter from Eq. (7). The standard deviation in this experiment is generated from a Gaussian distribution, scaled by a 1% relative error, and combined with the original data to simulate realistic measurement errors.
@@ III-C Experiment III:
In this work, we proposed approximating the magnetotelluric inverse operator using a neural network with additive layers. This approach enables the derivation of 1D resistivity models from impedance data. Our scheme is unsupervised and relies solely on a differentiable modeling operator constructed with tensor operations, allowing its integration as a forward operator in the neural network’s output layer. Using synthetic scenarios and field data, we compared our method with SimPEG and ZondMT1D. The results demonstrated that our approach better fits the observed data and produces subsurface models with structures similar to those obtained by the compared methods. Quantitatively, our method achieved the best performance metrics.

### v3

@@ Abstract
The global demand for unconventional energy sources such as geothermal energy and white hydrogen requires new exploration techniques for precise subsurface structure characterization and potential reservoir identification. The Magnetotelluric (MT) method is crucial for these tasks, providing critical information on the distribution of subsurface electrical resistivity at depths ranging from hundreds to thousands of meters. However, traditional iterative algorithm-based inversion methods require the adjustment of multiple parameters, demanding time-consuming and exhaustive tuning processes to achieve proper cost function minimization. Recent advances have incorporated deep learning algorithms for MT inversion, primarily based on supervised learning, and large labeled datasets are needed for training. This work utilizes TensorFlow operations to create a differentiable forward MT operator, leveraging its automatic differentiation capability. Moreover, instead of solving for the subsurface model directly, as classical algorithms perform, this paper presents a new deep unsupervised inversion algorithm guided by physics to estimate 1D MT models. Instead of using datasets with the observed data and their respective model as labels during training, our method employs a differentiable modeling operator that physically guides the cost function minimization, making the proposed method solely dependent on observed data. Therefore, the optimization algorithm updates the network weights to minimize the data misfit. We test the proposed method with field and synthetic data at different acquisition frequencies, demonstrating that the resistivity models obtained are more accurate than those calculated using other techniques. Our implementation is available at https://github.com/PAULGOYES/MT_guided1DInversion.git.
Index Terms:
1D magnetotelluric inversion, geophysical parameter estimation, deep unsupervised inversion, electromagnetic method, differentiable physical operator.I Introduction
@@ III-A Experiment I:
We evaluated the proposed method on a simple three-layer model with theoretical thicknesses of 2500 m for each layer and the depth to the top of the half-space layer at 5000 m. The MT acquisition was simulated with a frequency range from 10-3 Hz to 102 Hz. For inversion, we used a ten-layer model with a thickness of 555 m for each layer. In this experiment, we compared the proposed method with the solution of (6) i.e., –norm, which was initialized with a 100 m constant model. Note that for both cases, we used the same forward MT operator. We used a standard deviation in this experiment.
Fig. 2 shows the convergence curve for the cost function minimization. The –norm exhibits a smooth curve stabilizing after iteration 200, reaching a minimum error of approximately 0.35. On the other hand, the proposed method, which includes the inverse neural operator, displays a steep curve in the first five iterations and stabilizes after iteration 125 with small fluctuations around approximately 0.15. This indicates that the proposed method achieves faster convergence of the data misfit for the inverse problem. Although both curves converged, Fig. 3a shows that the –norm formulation exhibits a mismatch, particularly at the lowest frequencies, whereas the proposed method fits all observations across all frequencies. This translates to differences in the inverted models in Fig. 3b, mainly in the last layer (i.e., at a depth of 5 km), where the proposed method shows a good fit. The computation times are 85.543 s and 81.346 s for the proposed and the –norm formulation, respectively.
@@ III-B Experiment II:
We tested the inversion algorithm’s capability on data where subsurface layers have variable thicknesses to simulate a realistic subsurface scenario. Our results were compared with the SimPEG inversion framework [9, 21], which incorporates modeling operators based on Maxwell’s solutions to simulate electromagnetic fields with plane waves and a Gauss-Newton approach that employs the conjugate gradient solver [22]. SimPEG was initialized with a homogeneous resistivity model of 100 m. We used 20 layers with a maximum depth of 15,473 meters for both inversion methods. The proposed method used epochs. In this experiment, we found the best parameter from Eq. (7). The standard deviation in this experiment is generated from a Gaussian distribution, scaled by a 1% relative error, and combined with the original data to simulate realistic measurement errors.
@@ III-C Experiment III:
In this work, we proposed approximating the magnetotelluric inverse operator using a neural network with additive layers. This approach enables the derivation of 1D resistivity models from impedance data. Our scheme is unsupervised and relies solely on a differentiable modeling operator constructed with tensor operations, allowing its integration as a forward operator in the neural network’s output layer. We compared our method with SimPEG and ZondMT1D using synthetic scenarios and field data. The results demonstrated that our approach better fits the observed data and produces subsurface models with structures similar to those obtained by the compared methods. Quantitatively, our method achieved the best performance metrics.